# Joint Channel and Data Estimation for Multiuser Extremely Large-Scale MIMO Systems

### v2

@@ V-C1 Update $\bm{\pi}^{q,x}_{c,u,k}$
where $\bar{p}^{x}_{c,u,k}(\mathbf{y}_{c,k}|x_{u,k})$ is the conditional probability distribution defined in (35) in the top of next page along with $v^{e}_{n_{c},u,k}(e_{n_{c},u})$, and $\tilde{v}^{e}_{c,u,k}(\mathbf{e}_{c,u})$, with $\bar{C}^{x}_{c,u,k}$ being the normalizing constant and
@@
$\mathbf{\Omega}^{x}_{c,k}$ $=\sum_{u^{\prime}\in\mathcal{U}}\Big{\{}\xi^{c,x}_{c,u^{\prime},k}\hat{\mathbf{h}}^{v}_{c,u^{\prime},k}\hat{\mathbf{h}}^{v\mathrm{H}}_{c,u^{\prime},k}$
@@
Unlike the conventional MRC-based detections [13, 14, 12], the LMMSE-based detection for each sub-array $c$ in (40a) can deal with the correlation between the leaked energy in the beam-domain owing to whitening operation by $(\mathbf{\Omega}_{c,k}^{x})^{-1}$.
@@ V-C2 Update $\bm{\pi}^{b,x}_{u,k}$
from which the associated mean and variance are given by
@@
$b^{x}_{u,k}(x_{u,k})$ $\propto\ g(x_{u,k}|\mathbf{Y}_{u,k})\Big{/}\prod_{c^{\prime}\in\mathcal{C}}q^{x}_{c^{\prime},u,k}(x_{u,k}),$ (51)
@@ V-F Reinforcement for the Model-Based Estimate
To further improve the convergence performance for the EP algorithm, we update the model-based estimate $\hat{\mathbf{S}}$ in each iteration. Using the estimated residual channel error $\hat{\mathbf{E}}^{(t-1)}\triangleq\mathbb{E}_{g^{(t-1)}(\mathbf{E},\mathbf{X}|\mathbf{Y})}[\mathbf{E}]$ at the $(t-1)$-th iteration, the channel estimate for the $u$-th UE can be reconstructed as
@@
with $g_{\theta}\!\in\!\{1,\ldots,\bar{G}_{\theta}\}$, $g_{r}\!\in\!\{1,\ldots,\bar{G}_{r}\}$. $\hat{\theta}_{u,l}^{(t-1)}$ and $\hat{r}_{u,l}^{(t-1)}$ are the angle and distance estimates at the $(t-1)$-th iteration, respectively, and $\sigma_{\theta}^{(t)}$ and $\sigma_{r}^{(t)}$ are, respectively, the range of angle and distance grids, where the initial values $\hat{\theta}_{u,l}^{(0)}$ and $\hat{r}_{u,l}^{(0)}$ are determined using the angle and distance estimates obtained by the initial channel estimation as shown in Algorithm 1.
@@
with $\hat{\mathbf{z}}_{u}^{(t)}$ denoting the path gain estimates, $\mathbf{A}(\hat{\bm{\theta}}_{u}^{(t)},\hat{\mathbf{r}}_{u}^{(t)})$ being the corresponding array responses, which can be computed by solving
@@
To summarize, the proposed algorithm is encapsulated in Algorithm 2, where a damping scheme [16] is introduced in line 6, 7, 12, and 18 to enhance convergence performance.
@@
1:Input: $\mathbf{Y},\ \mathbf{X}_{\mathrm{p}},\ \mathbf{H}^{\mathcal{S}}_{0},\ \{\hat{\bm{\theta}}_{u},\hat{\mathbf{r}}_{u}\}_{u\in\mathcal{U}},\ \{\hat{L}_{u}\}_{u=1}^{U},\ T,\ \bar{G}_{\theta},\ \bar{G}_{r},$ $\sigma_{\theta}^{(1)},\ \sigma_{\theta}^{(T)},\ \sigma_{r}^{(1)},\ \sigma_{r}^{(T)}$
@@
5:$\hat{\mathbf{S}}=\mathbf{D}_{N}\mathbf{H}_{0}^{\mathcal{S}}$ from Algorithm 1
6:$\hat{x}^{v}_{c,u,k_{\mathrm{d}}}\!=\hat{x}^{w}_{c,u,k_{\mathrm{d}}}\!=0$$\xi^{v,x}_{c,u,k_{\mathrm{d}}}\!=\xi^{w,x}_{c,u,k_{\mathrm{d}}}\!=E_{s}$, $\forall k_{\mathrm{d}}\in\mathcal{K}_{\mathrm{d}}$
7:$\hat{x}^{v}_{c,u,k_{\mathrm{p}}}\!\!\!\!=\hat{x}^{w}_{c,u,k_{\mathrm{p}}}\!\!\!\!=[\mathbf{X}_{\mathrm{p}}]_{u,k_{p}}$$\xi^{v,x}_{c,u,k_{\mathrm{p}}}\!\!\!\!=\xi^{w,x}_{c,u,k_{\mathrm{p}}}\!\!\!\!=0$, $\forall k_{\mathrm{p}}\in\mathcal{K}_{\mathrm{p}}$
8:$\hat{e}^{v}_{n_{c},u,k}=0$$\xi^{v,e}_{n_{c},u,k}=|[\mathbf{D}_{N}\mathbf{H}_{0}^{\mathcal{S}}]_{n_{c},u}|^{2}$, $\forall k\in\mathcal{K}$
@@
15:     Calculate $\hat{x}^{v}_{c,u,k},\ \hat{\xi}^{v,x}_{c,u,k}$ from (50) with damping
16:     Calculate $\hat{x}^{w}_{c,u,k},\ \hat{\xi}^{w,x}_{c,u,k}$ from (56c) with damping
@@
22:     Calculate $\hat{e}^{v}_{n_{c},u,k},\hat{\xi}^{v,e}_{n_{c},u,k}$ from (63a)-(63b) with damping
23:     Update channel estimate $\hat{\mathbf{h}}_{u}=\hat{\mathbf{s}}_{u}+\hat{\mathbf{e}}_{u}$ from (68)
24: // EM algorithm for hyper parameter learning
25:     Update $\sigma^{e}_{n_{c},u}$ from (67)
26: // Reinforcement for the model-based estimate
27:     Generate the grids $\tilde{\theta}_{u,l,g_{\theta}},\ \tilde{r}_{u,l,g_{r}}$ from (69a)-(69b)
28:     Design the dictionary $\tilde{\mathbf{A}}_{u}(\tilde{\bm{\theta}}_{u},\tilde{\mathbf{r}}_{u})$ from (70)
29:     Obtain $\hat{\bm{\theta}}_{u},\hat{\mathbf{r}}_{u},\hat{\mathbf{z}}_{u}$ by solving (V-F)
30:     Calculate $\hat{\mathbf{s}}_{u}$ from (72) with damping
31:end for
@@ VI Simulation Results
This section evaluates the performance of the proposed initial channel estimation and subsequent JCDE algorithms under the following setup. The carrier frequency is $100\ \mathrm{GHz}$, the number of BS antennas $N$ is $200$, the number of UEs $U$ is $50$, the modulation order $Q$ is $64$-QAM, and the length of pilots $K_{\mathrm{p}}$ and data $K_{\mathrm{d}}$ are $25$ and $100$, respectively. The non-orthogonal pilot $\mathbf{X}_{\mathrm{p}}\in\mathbb{C}^{50\times 25}$ is designed by the frame design method in [15]. The near-field channel is composed of $L_{u}=3$ paths, i.e., $1$ LoS path and $2$ NLoS paths, with a Rician $K$-factor of 10 dB. The total number of paths is $L\!=\!50\!\times\!3\!=\!150$, and the corresponding oversampling quantity used in Algorithm 1 is set to $\hat{L}=250$. The AoAs and distances are uniformly randomly generated in the range $\theta_{u,l}\in[-60^{\circ},60^{\circ}]$ and $r_{u,l}\in[1,10]$ m, respectively. The polar-domain dictionary $\tilde{\mathbf{A}}(\tilde{\bm{\theta}},\tilde{\mathbf{r}})$ in (IV-A) is designed with $G_{r}=7$, $G_{\theta}=395$ and desired coherence $\gamma_{\mathrm{d}}=0.6$ in [10]. The performance is evaluated by the normalized mean-squared error (NMSE) and bit error rate (BER) under various signal-to-noise ratio (SNR). NMSE and SNR are defined as $\mathrm{NMSE}(\bm{\mathbf{H}})\triangleq\mathbb{E}\left[{\|\bm{\mathbf{H}}-\hat{\bm{\mathbf{H}}}\|_{\mathrm{F}}^{2}}/{\|\bm{\mathbf{H}}\|_{\mathrm{F}}^{2}}\right]$, and $\mathrm{SNR}\triangleq\mathbb{E}\left[\|\mathbf{HX}\|_{\mathrm{F}}^{2}\right]/\mathbb{E}\left[\|\mathbf{N}\|_{\mathrm{F}}^{2}\right]$. In what follows, the initial channel estimation and JCDE performance are evaluated in Section VI-A and VI-B, respectively.
@@ VI-A Initial Channel Estimation Performance
$+\hat{L}^{2}(G_{r}G_{\theta}K_{\mathrm{p}}+NU+NK_{\mathrm{p}})+\hat{L}^{3}K_{\mathrm{p}}+\hat{L}^{4}\Big{)}$
@@ VI-B JCDE Performance
In this subsection, we evaluate the performance of the proposed JCDE algorithm. As for JCDE algorithm parameters, the damping factor is set to $0.5$, the number of iterations is $T=30$, the number of grids are $\bar{G}_{\theta}=5,\bar{G}_{r}=5$, the grid ranges are $\sigma_{\theta}^{(1)}=5^{\circ}$, $\sigma_{\theta}^{(T)}=0.1^{\circ}$, $\sigma_{r}^{(1)}=5\ \mathrm{m}$, and $\sigma_{r}^{(T)}=1\ \mathrm{m}$, respectively. The extremely large array with $N=200$ antennas is divided into $C=4$ sub-arrays with $N_{c}=50$ antennas per sub-array. For comparison, AoA-aided BiGaBP [12] are employed as a benchmark, which is a state-of-the-art JCDE algorithm. Besides, we consider an ideal Genie-aided case with perfect knowledge of CSI or data corresponding to the lower bound of the proposed method.
@@ VI-B1 JCDE Performance with Initial Channel Estimation
Moreover, the proposed JCDE algorithm demonstrates a notable BER performance compared to the state-of-the-art AoA-aided BiGaBP [12]. The performance improvement can be attributed to two primary factors. The first factor is that the proposed algorithm can leverage the near-field model-based estimation described in Section V-F, whereas BiGaBP relies on the far-field assumption. The second factor is that the proposed sub-array-wise LMMSE-based detection in (40a) is capable of addressing the correlation between the leaked energy in the beam-domain, whereas BiGaBP is incapable of doing so because of its MRC. To reveal the aforementioned two factors, in Section VI-B2, we show the convergence analysis with and without near-field model information. Besides, we evaluate in Section VI-B3 the proposed sub-array-wise LMMSE-based detection performance and its complexity across various numbers of sub-arrays $C$.
@@ VI-B2 Convergence Analysis
To clarify the advantages gained by leveraging the near-field model structure, we evaluate the proposed JCDE algorithm with and without the model-based estimation process explained in Section V-F. Fig. 6 illustrates the BER and NMSE convergence behavior with respect to the number of algorithmic iterations. In the figure, the red triangle marker corresponds to the proposed JCDE algorithm without the model-based estimate, i.e., $\hat{\mathbf{S}}^{(t)}=\mathbf{0}$, where the prior distribution is designed i.i.d. for each element of $\mathbf{H}$ instead of $\mathbf{E}$, akin to [13, 14]. The green square marker corresponds to the proposed JCDE algorithm with the initial model-based estimate but without updating in iterations, i.e., $\hat{\mathbf{S}}^{(t)}=\hat{\mathbf{S}}^{(0)}$. Comparing the red triangle maker and green square marker, we can verify the performance improvement stemming from the use of the near-filed model through the decomposition of $\mathbf{H}$ into $\hat{\mathbf{S}}$ and $\mathbf{E}$ as written in (15). Furthermore, in comparison to the proposed algorithm with adaptive update, it can be seen that the adaptive updating of the model-based estimate $\hat{\mathbf{S}}^{(t)}$ enhances the BER and NMSE performance by further exploiting the near-field model.
@@
VI-B3 Performance Against the Number of Sub-arrays

### v3

@@ V-C1 Update $\bm{\pi}^{q,x}_{c,u,k}$
where $\bar{p}^{x}_{c,u,k}(\mathbf{y}_{c,k}|x_{u,k})$ is the conditional probability distribution defined in (35) at the top of next page along with $v^{e}_{n_{c},u,k}(e_{n_{c},u})$, and $\tilde{v}^{e}_{c,u,k}(\mathbf{e}_{c,u})$, with $\bar{C}^{x}_{c,u,k}$ being the normalizing constant and
@@
$\mathbf{\Omega}^{x}_{c,k}$ $=\sum_{u^{\prime}\in\mathcal{U}}\Big{\{}\xi^{v,x}_{c,u^{\prime},k}\hat{\mathbf{h}}^{v}_{c,u^{\prime},k}\hat{\mathbf{h}}^{v\mathrm{H}}_{c,u^{\prime},k}$
@@
Unlike the conventional MRC-based detections [12, 13, 11], the LMMSE-based detection for each sub-array in (40a) can deal with the correlation between the leaked energy in the beam-domain owing to the whitening operation by $(\mathbf{\Omega}_{c,k}^{x})^{-1}$. The correlation within the $c$-th sub-array, corresponding to the off-diagonal elements of $\mathbf{\Omega}^{x}_{c,k}$, increases as the leaked energy increases because the off-diagonal elements is computed by the weighted summation of the outer product of channel replicas $\hat{\mathbf{h}}^{v}_{c,u^{\prime},k}\hat{\mathbf{h}}^{v\mathrm{H}}_{c,u^{\prime},k}$ in (38). As illustrated in Fig. 2, the energy leakage effects become prominent as the distance between the BS and the UE decrease due to the near-field effects. Therefore, when the number of antennas per sub-array $N_{c}$ is small, including conventional MRC-based detections ($N_{c}=1$), each sub-array cannot address the correlation caused by the energy leakage. The number of antennas per sub-array $N_{c}$ should be selected by considering the trade-off between data detection performance and its computational complexity, which are evaluated through numerical simulations in Section VI.
@@ V-C2 Update $\bm{\pi}^{b,x}_{u,k}$
where the associated mean and variance are given by
@@
$b^{x}_{u,k}(x_{u,k})\$$\propto\ g(x_{u,k}|\mathbf{Y}_{u,k})\Big{/}\prod_{c^{\prime}\in\mathcal{C}}q^{x}_{c^{\prime},u,k}(x_{u,k}),$ (51)
@@ V-F Reinforcement for the Model-Based Estimate
To further improve the convergence performance for the EP algorithm, we update the model-based estimate $\hat{\mathbf{S}}$ in each iteration. Using the estimated residual channel error $\hat{\mathbf{E}}^{(t-1)}\triangleq\mathbb{E}_{g^{(t-1)}(\mathbf{E},\mathbf{X}|\mathbf{Y})}[\mathbf{E}]$ at the (t1)(t-1)-th iteration, the channel estimate for the $u$-th UE can be reconstructed as
@@
with $g_{\theta}\!\in\!\{1,\ldots,\bar{G}_{\theta}\}$, $g_{r}\!\in\!\{1,\ldots,\bar{G}_{r}\}$. $\hat{\theta}_{u,l}^{(t-1)}$ and $\hat{r}_{u,l}^{(t-1)}$ are the angle and distance estimates at the (t1)(t-1)-th iteration, respectively, and $\sigma_{\theta}^{(t)}$ and $\sigma_{r}^{(t)}$ are, respectively, the range of angle and distance grids, where the initial values $\hat{\theta}_{u,l}^{(0)}$ and $\hat{r}_{u,l}^{(0)}$ are determined using the angle and distance estimates obtained by the initial channel estimation as shown in Algorithm 1.
@@
with $\hat{\mathbf{z}}_{u}^{(t)}$ denoting the path gain estimates, $\mathbf{A}(\hat{\bm{\theta}}_{u}^{(t)},\hat{\mathbf{r}}_{u}^{(t)})$ being the array responses, which can be computed by solving
@@
Solving the problem (V-F) yields the path gain estimate $\hat{\mathbf{z}}_{u}^{(t)}$, angle estimate $\hat{\bm{\theta}}_{u}^{(t)}$ and the distance estimate $\hat{\mathbf{r}}_{u}^{(t)}$, which correspond to the non-zero elements of $\tilde{\mathbf{z}}_{u}$.
V-G Algorithm Description
To illustrate the overall procedure of the proposed JCDE algorithm, the pseudocode is provided in Algorithm 2, and the algorithmic flow is depicted in Fig. 4. As shown in the figure, the algorithm mainly consists of three parts: 1) data estimation, 2) residual error estimation, and 3) reinforcement for the model-based estimate.
In the data estimation, the mean and variance of the tentative data estimates $\{\hat{x}^{q}_{u,k},\xi^{q,x}_{u,k}\}$ are generated through the Soft-IC in (38a)-(38), sub-array-wise LMMSE filtering in (40a)-(40b), and combining processes (46), using the received signal $\mathbf{Y}$ and the soft replicas $\{\hat{x}^{v}_{c,u,k},\xi^{v,x}_{c,u,k}\}$ and $\{\hat{e}^{v}_{n_{c},u,k},\xi^{v,e}_{n_{c},u,k}\}$. Then, the tentative estimates $\{\hat{x}^{q}_{u,k},\xi^{q,x}_{u,k}\}$ are updated to $\{\hat{x}_{u,k},\xi^{x}_{u,k}\}$ through the denoising process in (48a)-(48b) that exploits the prior $p(x_{u,k})$ in (18) based on the QAM constellation. For the Soft-IC process in the next iteration, the soft replicas $\{\hat{x}^{v}_{c,u,k},\xi^{v,x}_{c,u,k}\}$ and $\{\hat{x}^{w}_{c,u,k},\xi^{w,x}_{c,u,k}\}$ are updated by suppressing the self-noise feedback in (50) and (56c).
In the residual error estimation, similar to the data estimation, the mean and variance of the tentative residual error estimates $\{\hat{e}^{q}_{n_{c},u},\xi^{q,e}_{n_{c},u}\}$ are generated through the Soft-IC in (56a)-(56) and MRC filtering processes in (61), using the received signal $\mathbf{Y}$ and the soft replicas $\{\hat{x}^{w}_{c,u,k},\xi^{w,x}_{c,u,k}\}$ and $\{\hat{e}^{v}_{n_{c},u,k},\xi^{v,e}_{n_{c},u,k}\}$. Then, the tentative estimates $\{\hat{e}^{q}_{n_{c},u},\xi^{q,e}_{n_{c},u}\}$ are updated to $\{\hat{e}_{n_{c},u},\xi^{e}_{n_{c},u}\}$ through the denoising process in (60) that exploits the prior $p(e_{n_{c},u};\mathbf{\Theta})$ in (19) reflecting the sparsity of the residual errors. For the Soft-IC process in the next iteration, the soft replicas $\{\hat{e}_{n_{c},u,k}^{v},\xi^{v,e}_{n_{c},u,k}\}$ are updated by suppressing the self-noise feedback in (63a)-(63b). The hyperparameter $\mathbf{\Theta}$ incorporated in the prior $p(e_{n_{c},u};\mathbf{\Theta})$ is updated through the M-step of the EM algorithm in (67). Based on the updated residual error estimate $\hat{\mathbf{e}}_{u}$, the channel estimate is reconstructed as $\hat{\mathbf{h}}_{u}=\hat{\mathbf{s}}_{u}+\hat{\mathbf{e}}_{u}$, and the model-based estimate $\hat{\mathbf{s}}_{u}$ is subsequently refined by solving the optimization problem (V-F) via OMP [23].
After repeating the above procedure for a fixed number of iterations $T$, $\hat{\mathbf{X}}$ and $\hat{\mathbf{H}}$ are adopted as the estimate of the data and channel. To enhance convergence performance, a damping scheme [15] is introduced in line 6, 7, 12, and 19 of Algorithm 2. The damping factor in the proposed method is determined through several simulations by considering the trade-off between estimation performance and convergence rate as in [10, 26]. Note that the proposed initial channel estimation in Algorithm 1 and the JCDE in Algorithm 2 does not require distinguishing between the LoS and NLoS paths.
@@ V-G Algorithm Description
1:Input: $\mathbf{Y},\ \mathbf{X}_{\mathrm{p}},\ \hat{\mathbf{H}}^{\mathcal{S}}_{0},\ \{\hat{\bm{\theta}}_{u},\hat{\mathbf{r}}_{u}\}_{u\in\mathcal{U}},\ \{\hat{L}_{u}\}_{u=1}^{U},\ T,\ \bar{G}_{\theta},\ \bar{G}_{r},$ $\sigma_{\theta}^{(1)},\ \sigma_{\theta}^{(T)},\ \sigma_{r}^{(1)},\ \sigma_{r}^{(T)}$
@@
[…]
@@
15:  Calculate $\hat{x}^{v}_{c,u,k},\ \xi^{v,x}_{c,u,k}$ from (50) with damping
16:  Calculate $\hat{x}^{w}_{c,u,k},\ \xi^{w,x}_{c,u,k}$ from (56c) with damping
@@
22:  Calculate $\hat{e}^{v}_{n_{c},u,k},\xi^{v,e}_{n_{c},u,k}$ from (63a)-(63b) with
23:   damping
24:  Update channel estimate $\hat{\mathbf{h}}_{u}=\hat{\mathbf{s}}_{u}+\hat{\mathbf{e}}_{u}$ from (68)
25: // EM algorithm for hyperparameter learning
26:  Update $\sigma^{e}_{n_{c},u}$ from (67)
27: // Reinforcement for the model-based estimate
28:  Generate the grids $\tilde{\theta}_{u,l,g_{\theta}},\ \tilde{r}_{u,l,g_{r}}$ from (69a)-(69b)
29:  Design the dictionary $\tilde{\mathbf{A}}_{u}(\tilde{\bm{\theta}}_{u},\tilde{\mathbf{r}}_{u})$ from (70)
30:  Obtain $\hat{\bm{\theta}}_{u},\hat{\mathbf{r}}_{u},\hat{\mathbf{z}}_{u}$ by solving (V-F)
31:  Calculate $\hat{\mathbf{s}}_{u}$ from (72) with damping
32:end for
@@ VI Simulation Results
[…]
@@ VI-A Initial Channel Estimation Performance
$\hat{L}^{2}(G_{r}G_{\theta}K_{\mathrm{p}}+NU+NK_{\mathrm{p}})+\hat{L}^{3}K_{\mathrm{p}}+\hat{L}^{4}\Big{)}$
@@ VI-B JCDE Performance
In this subsection, we evaluate the proposed JCDE algorithm. As for JCDE algorithm parameters, the damping factor is set to $0.5$, the number of iterations is $T=30$, the number of grids are $\bar{G}_{\theta}=5,\bar{G}_{r}=5$, the grid ranges are $\sigma_{\theta}^{(1)}=5^{\circ}$, $\sigma_{\theta}^{(T)}=0.1^{\circ}$, $\sigma_{r}^{(1)}=5\ \mathrm{m}$, and $\sigma_{r}^{(T)}=1\ \mathrm{m}$, respectively. To consider the simplicity of implementation and the complexity of the LMMSE-based filter in (40a), the large array with $N$ antennas is evenly divided into $C$ sub-arrays with $N_{c}=N/C,\ \forall c\in\mathcal{C}$ antennas. Since the complexity order of the LMMSE-based filter is $\mathcal{O}\left(K_{\mathrm{d}}\sum_{c=1}^{C}N_{c}^{3}+UK_{\mathrm{d}}\sum_{c=1}^{C}N_{c}^{2}\right)$, even sub-array partitioning is optimal in terms of minimizing computational complexity given $C$. Unless otherwise specified, these parameters are set to $N=200,\ C=4,\ N_{c}=50$. For comparison, AoA-aided BiGaBP [11] are employed as a benchmark, which is a state-of-the-art JCDE algorithm. Besides, we consider a Genie-aided case with perfect CSI or data corresponding to the lower bound of the proposed method.
@@ VI-B1 JCDE Performance with Initial Channel Estimation
Moreover, the proposed JCDE algorithm demonstrates a notable BER performance compared to the state-of-the-art AoA-aided BiGaBP [11]. The performance improvement can be attributed to two primary factors. The first factor is that the proposed algorithm can leverage the near-field model-based estimation described in Section V-F, whereas BiGaBP relies on the far-field assumption. The second factor is that the proposed sub-array-wise LMMSE-based detection in (40a) is capable of addressing the correlation between the leaked energy in the beam-domain, whereas BiGaBP is incapable of doing so because of its MRC. To reveal the aforementioned two factors, in Section VI-B2, we show the convergence analysis with and without near-field model information. Besides, we evaluate in Section VI-B3 the performance of the sub-array-wise LMMSE-based detection and its complexity across various numbers of antennas per sub-array $N_{c}$.
@@ VI-B2 Convergence Analysis
To clarify the advantages gained by leveraging the near-field structure, we evaluate the proposed JCDE algorithm with and without the model-based estimation process explained in Section V-F. Fig. 7 illustrates the BER and NMSE convergence behavior with respect to the number of algorithmic iterations. In the figure, the red triangle marker corresponds to the proposed JCDE algorithm without the model-based estimate, i.e., $\hat{\mathbf{S}}^{(t)}=\mathbf{0}$, where the prior distribution is designed i.i.d. for each element of $\mathbf{H}$ instead of $\mathbf{E}$, akin to [12, 13]. The green square marker corresponds to the proposed JCDE algorithm with the initial model-based estimate but without updating in iterations, i.e., $\hat{\mathbf{S}}^{(t)}=\hat{\mathbf{S}}^{(0)}$. Comparing the red triangle maker and green square marker, we can verify the performance improvement owing to the use of the near-filed model through the decomposition of $\mathbf{H}$ into $\hat{\mathbf{S}}$ and $\mathbf{E}$ in (15). Furthermore, in comparison to the proposed algorithm with adaptive update, it can be seen that the adaptive updating of the model-based estimate $\hat{\mathbf{S}}^{(t)}$ enhances the BER and NMSE by exploiting the near-field model.
@@
VI-B3 Performance Against the Number of Antennas per Sub-array